OpenStudy (anonymous):

Ty recently bought a car for $42,000 His research shows the car will depreciate by an average of $800 per year If x represents the # of yrs ty owns the car which of the following equations represents the value of the car after x yrs

OpenStudy (anonymous):

y = 800x - 42,000 y = 42,000x - 800 y = 42,000 - 800x y = 42,000 + 800x

OpenStudy (anonymous):

3rd one

OpenStudy (anonymous):

definetly 3rd

OpenStudy (bibby):

it has an initial starting price of 42,000 and loses 800 per year. meaning on the first year, it loses 800. on the 2nd 1600. C is right